# OmniSVG: A Unified Scalable Vector Graphics Generation Model

### v2

@@ 6 Conclusions
Figure 9: Limitation of OmniSVG on Image-to-SVG Task. OmniSVG can successfully generate vector style images, while fail to fit natural images.
Conclusions. We introduce OmniSVG, a unified framework for multimodal SVG generation that leverages pre-trained Vision-Language Models (VLMs). By parameterizing SVG commands and coordinates as discrete tokens, OmniSVG efficiently decouples structural logic from geometry, addressing issues like ”coordinate hallucination” while maintaining design expressiveness. Our method outperforms existing approaches in both quality and efficiency, offering high-quality, editable SVG across various design domains. Additionally, we proposed MMSVG-2M, a large-scale multimodal dataset with two million annotated SVG assets and a standardized evaluation protocol. Extensive experiments show that OmniSVG surpasses prior SVG generation methods in various conditional generation tasks, highlighting its potential for integration into professional SVG design workflows.
Limitations and Future Work. During inference, OmniSVG generates tens of thousands of tokens for complex samples, which inevitably leads to a considerable generation time. OmniSVG is only bounded by vector style image prompt and fails on natural images as illustrated in Fig. 9. As for future work, recent endeavors on multi-token prediction [15, 2] and KV-cache compression [68, 3] provide a promising way to save the generation cost. Additionally, the auto-regressive nature of OmniSVG also unlocks future opportunities for in-context learning [67, 69, 46], chain-of-thought reasoning [54, 16], and multi-turn interleaved generation [19, 30], thereby providing a more dexterous user control. Moreover, more complicated tasks like stylization [65, 70] is also worth exploring in the future.
@@ A.1 Samples of MMSVG-2M Dataset
We visualize samples of our MMSVG-2M dataset in Fig. 10. In our MMSVG-2M dataset, 55% of the SVG samples belongs to the MMSVG-Icon, 25% belongs to the MMSVG-Illustration, and the rest 20% belongs to the MMSVG-Character. Among the SVG samples within the MMSVG-Character category, half of them comes from Freepik, while another half is generated by our data creation pipeline. We also collect image-SVG pairs for the character-reference SVG generation tasks during the generation process.
@@
A.2 SVG-Image-Text Pairs Construction
Our MMSVG-2M dataset comprises two million SVG samples with the corresponding rasterized images. Following the recent dataset construction pipeline [28], we generate captions on the rasterized images with BLIP-2 [26], thereby providing textual descriptions that enable us to fine-tune our model to follow these instructions. We use CairoSVG [23] for rasterization and remove samples that produced completely white images.
@@ A.2 SVG-Image-Text Pairs Construction
Annotation. We employ an off-the-shelf VLM, specifically BLIP-2 [26], to generate SVG captions with the prompt below. To reduce hallucinations, we drop the samples with CLIP scores less than 30. We also visualize the distribution annotated keywords of MMSVG-2M dataset in Fig. 12 with word cloud format. And the instruction template for annotation is shown in Tab. 8.
@@
Instruction templates. MMSVGBench provides three tasks, including text-to-SVG task, image-to-SVG task and character-reference SVG generation task. Each task needs different instruction templates. For the text and image conditioning SVG generation, we provide the input text or image with VLM architecture. For character-reference SVG generation, we provide the natural charecter reference image and the original image with the VLM architecture. The list of instruction templates for different tasks are shown in Tab. 8.
Employed BLIP2 for SVG Captioning: You are a helpful assistant. Your task is to describe this image in a single sentence, including the object, its color, and its overall arrangement. For example: “Yellow cheers with glasses of alcohol drinks.” / “Heart emojis represent love on Valentine’s Day.” Text-to-SVG: You are a helpful SVG Generation assistant, designed to generate SVG. We provide the text description as input, generate SVG based on the text. Image-to-SVG: You are a helpful SVG Generation assistant, designed to generate SVG. We provide an image as input, generate SVG for this image. Character-Reference SVG Generation: You are a helpful SVG Generation assistant, designed to generate SVG. We provide two image as input, the second image is the character reference image of the first image, generate the character reference SVG based on these two input images.
Table 8: The list of instructions for different tasks, including annotation, text-to-SVG, image-to-SVG and character-reference SVG generation.
A.3 Character-SVG Pairs Construction
As illustrated in the Fig. 1, part of our proposed MMSVG-2M-Character subset is constructed using a generative pipeline. As shown in the pipeline diagram in Fig. 2, we employ a FLUX [24]-based generative model enhanced with a vector-style LoRA to enable the generation of SVG-style data. For image-based conditioning, we adopt FLUX-Redux [25], which injects image features via a SigLIP encoder and projects them into image embeddings. These embeddings are then concatenated with the text tokens as conditioning inputs for FLUX [24]. However, in practice, the original Redux [25] conditioning proves to be overly strong. To address this, we adopt a community-implemented variant of Redux that downsamples the image embeddings in 2D space. As observed in our experiments shown in Fig. 11, a downsampling factor between 2×2$\times$2 × and 3×3$\times$3 × yields the most reasonable SVG-style character references. Finally, we employ VTracer [12] to perform near-instant vectorization of the generated images. To construct the MMSVG-2M-Character subset, we first filter $103k$ character instances from the Danbooru [13] dataset and apply the aforementioned pipeline with motion and expression keywords like previous works [9, 8, 35, 64]. We compare the raw FLUX [24] outputs and their vectorized counterparts, retaining only those samples with PSNR and SSIM scores above a certain threshold as valid data.
Appendix B More details of the baselines
B.1 Text-to-SVG Task
@@
SVGDreamer [59] uses a semantic-driven image vectorization (SIVE) process to separate foreground objects and background, improving editability. The SIVE process utilizes attention-based primitive control and an attention-mask loss function to manipulate individual elements effectively. To address issues in existing text-to-SVG generation methods, the proposed Vectorized Particle-based Score Distillation (VPSD) approach models SVGs as distributions of control points and colors, improving shape, color diversity, and convergence speed.
@@
VectorFusion [21] leverages a text-conditioned diffusion model trained on pixel representations to generate SVG exportable vector graphics without needing large captioned SVG datasets. By optimizing a differentiable vector graphics rasterizer, it distills semantic knowledge from a pretrained diffusion model and uses Score Distillation Sampling to generate an SVG consistent with a caption. Experiments show that VectorFusion improves both quality and fidelity, offering a variety of styles such as pixel art and sketches.
@@
Chat2SVG [56] proposes a hybrid framework that combines the strengths of Large Language Models (LLMs) and image diffusion models for text-to-SVG generation. The approach first uses an LLM to create semantically meaningful SVG templates from basic geometric primitives. A dual-stage optimization pipeline, guided by image diffusion models, refines paths in latent space and adjusts point coordinates to enhance geometric complexity.
@@
IconShop [55] uses a transformer-based architecture to encode path commands and learn to model SVG path sequences autoregressively. It has shown excellent results in simplified icon scenarios and provides a good solution to Text-to-SVG generation by extending the FIGR-8-SVG dataset with captions. We have access to their dataset and original splits and have trained our model on that data using a pre-trained checkpoint (trained on OmniVG dataset). We have extracted the results from IconShop and included them here to compare our method.
@@
LLM4SVG [58] is a framework that leverages Large Language Models (LLMs) to understand and generate Scalable Vector Graphics (SVGs). It employs a structured SVG encoding approach, utilizing learnable semantic tokens to accurately represent SVG components and their properties. This design enables LLMs to produce SVGs that are both semantically aligned with textual descriptions and visually coherent. However, LLM4SVG also has a maximum token length of 2048, limiting its ability to generate highly complex SVGs that require longer sequences.
B.2 Image-to-SVG Task
@@
LIVE (Layer-wise Image Vectorization) [33] is a method for progressively generating SVGs that closely fit a given raster image by recursively adding and optimizing closed vector paths. Using a differentiable renderer (based on DiffVG [27]), LIVE enables direct optimization of paths under raster image supervision while controlling shape complexity by adjusting the number of path segments. It introduces component-wise path initialization, identifying key visual components to ensure efficient topology extraction and minimize redundant shapes.
@@
DiffVG [27] is a landmark in vector graphics research, pioneering deep learning-based methods with the first differentiable vector graphics rasterization pipeline. By leveraging a combination of anti-aliasing techniques and gradient-based optimization, DiffVG ensures differentiability. Unlike methods relying on non-differentiable curve-to-mesh conversions, DiffVG employs a forward-backward rasterization process, where the forward pass generates antialiased images and the backward pass computes gradients with respect to vector graphic parameters.
@@
StarVector [41] works directly in the SVG code space, leveraging visual understanding to apply accurate SVG primitives. StarVector employs a transformer-based architecture that integrates an image encoder with a language model, enabling it to process visual inputs and produce precise SVG code. StarVector effectively handles diverse SVG types, including icons, logos, and complex diagrams, demonstrating robust generalization across various vectorization tasks. However, with a 16k token context window, StarVector may struggle to process highly complex SVGs that require longer sequences.

### v3

@@ 6 Conclusions
Conclusions. We introduce OmniSVG, a unified framework for multimodal SVG generation that leverages pre-trained Vision-Language Models (VLMs). By parameterizing SVG commands and coordinates as discrete tokens, OmniSVG efficiently decouples structural logic from geometry, addressing issues like "coordinate hallucination" while maintaining design expressiveness. Our method outperforms existing approaches in both quality and efficiency, offering high-quality, editable SVG across various design domains. Additionally, we proposed MMSVG-2M, a large-scale multimodal dataset with two million annotated SVG assets and a standardized evaluation protocol. Extensive experiments show that OmniSVG surpasses prior SVG generation methods in various conditional generation tasks, highlighting its potential for integration into professional SVG design workflows.
Limitations and Future Work. During inference, OmniSVG generates tens of thousands of tokens for complex samples, which inevitably leads to a considerable generation time. OmniSVG is only bounded by vector style image prompt and fails on natural images. As for future work, recent endeavors on multi-token prediction gloeckle2024better; cai2023medusa and KV-cache compression zhou2024dynamickv; cai2024pyramidkv provide a promising way to save the generation cost. Additionally, the auto-regressive nature of OmniSVG also unlocks future opportunities for in-context learning zhang2023makes; zhou2024visual; sun2024generative, chain-of-thought reasoning wei2022chain; guo2025deepseek, and multi-turn interleaved generation huang2024dialoggen; liu2024mmdu, thereby providing a more precise user control.
Acknowledgements
This work is in part supported by National Key R&D Program of China (Grant No. 2022ZD0160103), National Natural Science Foundation of China (Grant No. 62276067), and National Natural Science Foundation of China (Grant No. 62472104).
The computations in this research were performed using the CFFF platform of Fudan University.
@@ A.1 Samples of MMSVG-2M Dataset
We visualize samples of our MMSVG-2M dataset in Fig.˜8. In our MMSVG-2M dataset, 55% of the SVG samples belongs to the MMSVG-Icon, 25% belongs to the MMSVG-Illustration, and the rest 20% belongs to the MMSVG-Character. Among the SVG samples within the MMSVG-Character category, half of them comes from Freepik, while another half is generated by our data creation pipeline. We also collect image-SVG pairs for the character-reference SVG generation tasks during the generation process.
Dataset Train Val Total Source Token Length
MMSVG-Icon 990k 110k 1,100k Iconfonthttps://www.iconfont.cn/ $2.2\text{k}\pm 0.9\text{k}$
MMSVG-Illustration 450k 50k 500k IconScout $8.1\text{k}\pm 3.3\text{k}$
MMSVG-Character 350k 50k 400k Freepik & generated $28\text{k}\pm 7.3\text{k}$
A.2 SVG-Image-Text Pairs Construction
Our MMSVG-2M dataset comprises two million SVG samples with the corresponding rasterized images. We generate captions on the rasterized images with BLIP-2 blip2, thereby providing textual descriptions that enable us to fine-tune our model to follow these instructions. We use CairoSVG cairosvg for rasterization and remove samples that produced completely white images.
Annotation. We employ an off-the-shelf VLM, specifically BLIP-2 blip2, to generate SVG captions with the prompt below. To reduce hallucinations, we drop the samples with CLIP scores less than 30. We also visualize the distribution annotated keywords of MMSVG-2M dataset in Fig. 10 with word cloud format. And the instruction template for annotation is shown in Tab.˜7.
Instructions for Different Tasks - Employed BLIP2 for SVG Captioning: You are a helpful assistant. Your task is to describe this image in a single sentence, including the object, its color, and its overall arrangement. For example: “Yellow cheers with glasses of alcohol drinks." / “Heart emojis represent love on Valentine’s Day." - Text-to-SVG: You are a helpful SVG Generation assistant, designed to generate SVG. We provide the text description as input, generate SVG based on the text. - Image-to-SVG: You are a helpful SVG Generation assistant, designed to generate SVG. We provide an image as input, generate SVG for this image. - Character-Reference SVG Generation: You are a helpful SVG Generation assistant, designed to generate SVG. We provide a natural image as input, please generate the simplified character SVG based on the reference input image.
Table 7: Instructions for Different Tasks. Instructions including annotation, text-to-SVG, image-to-SVG and character-reference SVG generation.
Instruction templates. MMSVGBench provides three tasks, including text-to-SVG task, image-to-SVG task and character-reference SVG generation task. Each task needs different instruction templates. For the text and image conditioning SVG generation, we provide the input text or image with VLM architecture. For character-reference SVG generation, we provide the natural charecter reference image and the original image with the VLM architecture. The list of instruction templates for different tasks are shown in Tab.˜7.
A.3 Character-SVG Pairs Construction
As illustrated in the Fig. 6, part of our proposed MMSVG-2M-Character subset is constructed using a generative pipeline. As shown in the pipeline diagram in Fig. 2, we employ a FLUX (flux2024)-based generative model enhanced with a vector-style LoRA to enable the generation of SVG-style data. For image-based conditioning, we adopt FLUX-Redux (redux), which injects image features via a SigLIP encoder and projects them into image embeddings. These embeddings are then concatenated with the text tokens as conditioning inputs for FLUX (flux2024). However, in practice, the original Redux (redux) conditioning proves to be overly strong. To address this, we adopt a community-implemented variant of Redux that downsamples the image embeddings in 2D space. As observed in our experiments shown in Fig. 9, a downsampling factor between $2\times$ and $3\times$ yields the most reasonable SVG-style character references. Finally, we employ VTracer (vtracer) to perform near-instant vectorization of the generated images. To construct the MMSVG-2M-Character subset, we first filter $103k$ character instances from the Danbooru (Danbooru2023) dataset and apply the aforementioned pipeline with motion and expression keywords like previous works cheng2023dna; cheng2022generalizable; pan2023renderme; yu2023monohuman. We compare the raw FLUX (flux2024) outputs and their vectorized counterparts, retaining only those samples with PSNR and SSIM scores above a certain threshold as valid data.
@@ A.3 Character-SVG Pairs Construction
Appendix B Additional Details
B.1 Scaling Up
To study the effectiveness of scaling up multimodal SVG generation, we scale up OmniSVG from 4B to 8B parameters. We present training perplexity in Fig. 11, where both models are trained from scratch on 250 billion tokens. We show that, as the size of the model grows, the model achieves a lower validation perplexity, indicating a higher probability of producing the validation data.
@@ B.1 Scaling Up
B.2 Implementation Details
We train our models in bfloat16 with the ZeRO-2 strategy zero for memory-efficient training. We also adopt the AdamW adamw optimizer with a learning rate decaying from $3\times 10^{-4}$ to $3\times 10^{-6}$ and a weight decay of $0.1$ to train our model. In practice, we load the pre-trained weights from the Qwen2.5-VL qwen2.5-VL model and initialize the SVG embeddings from scratch. Without further specification, we generate SVGs with the top-k and top-p sampling strategy with $k=50$ and $p=0.95$ for diversity.
(a) Training PPL for our models.
(b) Validation PPL for our models.
Appendix C Additional Results
As list in full comparisons in Tab.˜2, including all the baselines mentioned in Sec.˜5. For the text-to-SVG task, we compare our method with language-based (LLM-based) methods, including VectorFusion (vectorfusion), SVGDreamer (svgdreamer), Chat2SVG (chat2svg) and IconShop (iconshop). For image-to-SVG task, we compare our method with baseline methods across image vectorization and Multimodal Large Language Modeling approaches, including LIVE (live), DiffVG (diffvg), StarVector (starvector) and GPT-4o (GPT-4o) using the official implementations with the hyperparameters proposed by the authors, and apply their pre- and post-processing code as required. Specifically, for the text-to-SVG task, the optimization-based method SVGDreamer excels in enhancing editability by employing a semantic-driven image vectorization process that effectively separates foreground objects from the background, while failing to handle complex scenes. Another optimization-based work, VectorFusion, stands out for generating SVG-exportable vector graphics without relying on large captioned datasets. However, Vectorfusion is also unable to handle complex scenarios and diverse styles. The significant problem with these optimization-based works is that the optimization time is too long. Generating an SVG usually takes more than ten minutes, which is too expensive. For the LLM-based method, Chat2SVG integrates Large Language Models (LLMs) with image diffusion models to create semantically rich SVG templates. However, Chat2SVG still needs to optimize the output SVG script from LLM, which introduces increased computational complexity and poses challenges during model training. In comparison, IconShop utilizes a transformer-based architecture to autoregressively model SVG path sequences, demonstrating exceptional performance in simplified icon SVGs, which offers effective solutions for text-to-SVG generation. It can only generate black simple Icon SVGs.
For the image-to-SVG task, we compare our method with the image vectorization methods. LIVE allows progressive and efficient generation of SVGs, optimizing closed vector paths under raster image supervision with shape complexity control. However, LIVE needs to optimize for a long time when generating complex SVGs. DiffVG enables end-to-end differentiability in vector graphics rasterization, improving optimization through anti-aliasing and gradient-based methods while also is computationally expensive due to the complexity of the forward-backward rasterization process. Recently, the Multimodal Large Language Model (MLLM) based method StarVector leverages the visual understanding to apply accurate SVG primitive to the LLM architecture, which also can generate SVGs from both text and image inputs. However, it still fails to generate complex SVGs. Since Starvector (starvector) has not yet opened up its text-to-SVG model weights, our MMSVGBench does not evaluate Starvector’s text-to-SVG capabilities. MMSVG-Bench also evaluates our methods with VLM methods, GPT-4o, to conduct a comprehensive assessment. We compare our method with these baselines on our MMSVG-2M dataset, from simple MMSVG-Icon datset, a bit complex MMSVG-illustration dataset, to the very complex MMSVG-Character dataset.
Appendix D More details of the baselines
D.1 Text-to-SVG Task
@@
SVGDreamer svgdreamer uses a semantic-driven image vectorization (SIVE) process to separate foreground objects and background, improving editability. The SIVE process utilizes attention-based primitive control and an attention-mask loss function to manipulate individual elements effectively. To address issues in existing text-to-SVG generation methods, the proposed Vectorized Particle-based Score Distillation (VPSD) approach models SVGs as distributions of control points and colors, improving shape, color diversity, and convergence speed.
@@
VectorFusion vectorfusion leverages a text-conditioned diffusion model trained on pixel representations to generate SVG exportable vector graphics without needing large captioned SVG datasets. By optimizing a differentiable vector graphics rasterizer, it distills semantic knowledge from a pretrained diffusion model and uses Score Distillation Sampling to generate an SVG consistent with a caption. Experiments show that VectorFusion improves both quality and fidelity, offering a variety of styles such as pixel art and sketches.
@@
Chat2SVG chat2svg proposes a hybrid framework that combines the strengths of Large Language Models (LLMs) and image diffusion models for text-to-SVG generation. The approach first uses an LLM to create semantically meaningful SVG templates from basic geometric primitives. A dual-stage optimization pipeline, guided by image diffusion models, refines paths in latent space and adjusts point coordinates to enhance geometric complexity.
@@
IconShop iconshop uses a transformer-based architecture to encode path commands and learn to model SVG path sequences autoregressively. It has shown excellent results in simplified icon scenarios and provides a good solution to Text-to-SVG generation by extending the FIGR-8-SVG dataset with captions. We have access to their dataset and original splits and have trained our model on that data using a pre-trained checkpoint (trained on OmniVG dataset). We have extracted the results from IconShop and included them here to compare our method.
@@
LLM4SVG llm4svg is a framework that leverages Large Language Models (LLMs) to understand and generate Scalable Vector Graphics (SVGs). It employs a structured SVG encoding approach, utilizing learnable semantic tokens to accurately represent SVG components and their properties. This design enables LLMs to produce SVGs that are both semantically aligned with textual descriptions and visually coherent. However, LLM4SVG also has a maximum token length of 2048, limiting its ability to generate highly complex SVGs that require longer sequences.
D.2 Image-to-SVG Task
@@
LIVE (Layer-wise Image Vectorization) live is a method for progressively generating SVGs that closely fit a given raster image by recursively adding and optimizing closed vector paths. Using a differentiable renderer (based on DiffVG diffvg), LIVE enables direct optimization of paths under raster image supervision while controlling shape complexity by adjusting the number of path segments. It introduces component-wise path initialization, identifying key visual components to ensure efficient topology extraction and minimize redundant shapes.
@@
DiffVG diffvg is a landmark in vector graphics research, pioneering deep learning-based methods with the first differentiable vector graphics rasterization pipeline. By leveraging a combination of anti-aliasing techniques and gradient-based optimization, DiffVG ensures differentiability. Unlike methods relying on non-differentiable curve-to-mesh conversions, DiffVG employs a forward-backward rasterization process, where the forward pass generates antialiased images and the backward pass computes gradients with respect to vector graphic parameters.
@@
StarVector starvector works directly in the SVG code space, leveraging visual understanding to apply accurate SVG primitives. StarVector employs a transformer-based architecture that integrates an image encoder with a language model, enabling it to process visual inputs and produce precise SVG code. StarVector effectively handles diverse SVG types, including icons, logos, and complex diagrams, demonstrating robust generalization across various vectorization tasks. However, with a 16k token context window, StarVector may struggle to process highly complex SVGs that require longer sequences.
Vtracer vtracer is an image processing algorithm designed to convert raster images into SVGs. The algorithm follows a three-step pipeline, which involves the hierarchical clustering of images for vectorization. Initially, the pixels are transformed into paths, which are subsequently simplified into polygons. In the final step, these polygons are smoothed and approximated using a Bezier curve fitter.